OpenStudy (anonymous):

Find the average value of the function f(x)=1-|x| on the interval [-1,3]

OpenStudy (anonymous):

\[1/(b-a)\int\limits_{a}^{b}f(x)dx\]

OpenStudy (anonymous):

ans= -1/4

OpenStudy (anonymous):

hmm what did i do wrong here. 1/(3- (-1))= 1/4

OpenStudy (anonymous):

you haven't solved the integral.

OpenStudy (anonymous):

\[ 1/4*\int\limits_{-1}^{3} 1-|x| = 1/4* x-(x^2/2)\]

OpenStudy (anonymous):

solve the integral by graphing it. it's way easier.

OpenStudy (anonymous):